Functional Programming Column

Philip Wadler, University of Glasgow

                     wadler@dcs.glasgow.ac.uk
               http://www.dcs.glasgow.ac.uk/~wadler

This month’s functional programming column is short and sweet – an invitation to read or contribute to The Journal Of Functional Programming.

FACS readers are encouraged to submit to this column, as well as to JFP. Short papers on the following topics are especially welcome.

• Specifications written as functional programs.

• Proving properties of functional programs.

• Real-world applications of functional programming. (Please also submit

     these to our web page, http://www.dcs.gla.ac.uk/fp/realworld.html.)

• Tutorial summaries of relevant advances in FP.

Submissions are welcome at any time, electronically or by post.

Journal of Functional Programming: The state of play

Simon Peyton Jones and Philip Wadler

              http://www.dcs.gla.ac.uk/jfp

We’re delighted to announce the publication of a special issue of JFP 5(3) on applications of functional programming. Edited by Rinus Plasmeijer and Pieter Hartel, it contains six papers on applications ranging from photon transport to text retrieval, from spreadsheets to solid modelling.

The titles of the papers are given below. You can see their abstracts too on the above web page. You’ll also now find titles and abstracts for the whole of Volumes 4 and 5, and titles for Volumes 1–3.

We’d like to take this opportunity to encourage you to submit papers to JFP. At one point we had so many submissions that we built up rather a long print queue, but Cambridge University Press have agreed to move from 4 to 6 issues a year (starting in 1996), and to make them big issues, which has effectively solved the print-queue problem.

We offer a 12-week turnaround to a decision about submitted papers, and the delay from final manuscript to publication is now dropping fast (our target is 3 months).

Please submit papers on any aspect of functional programming to any of the Editors:

          Henk Barendregt  henk@cs.kun.nl
          Paul Hudak  hudak@cs.yale.edu
          John Hughes  rjmh@cs.chalmers.se
          Simon Peyton Jones simonpj@dcs.gla.ac.uk
          Philip Wadler  wadler@dcs.gla.ac.uk

You can find full details of the journal’s scope, editorial policy, Latex style files, and subscription information by following the web link above.

JFP Special Issue 5(3) (July 1995) on State-of-the-art applications of pure functional

programming

Editors: Pieter Hartel and Rinus Plasmeijer

• Comparing Id and Haskell in a Monte Carlo photon transport code, Jeffrey Hammes, Olaf Lubeck and Wim Bohm

• Funser: a functional server for textual information retrieval, Donald A Ziff, Stephen P Spackman and Keith Waclena

• Prototyping a parallel vision system in Standard ML, Greg Michaelson and Norman Scaife

• Implementing a functional spreadsheet in Clean Walter A.C.A.J.De Hoon, Luc M.W.J.Rutten and Marko C.J.D. Van Eekelen

• A polymorphic library for constructive solid geometry J.R. Davy and P.M. Dew

• Exploring the conformations of nucleic acids Marcel Turcotte, Guy Lapalme and Francois Major